Project

General

Profile

20151201 » History » Version 27

Boyd, Steven, 09 December 2015 07:36

1 11 Boyd, Steven
Commissioning and Ops Scheduling meeting : 8/12/2015
2 1 Boyd, Steven
3
To be held in CR10 in the ATLAS Building @ 15:30 pm 
4
5
Phone access :  ReadyTalk number 
6
US 866-740-1260
7
UK 0800-496-0576
8
Swiss 0800-70-53-51
9
Italy 800182592
10
Access Code: 2155402
11
12
h2. Agenda:
13
14 2 Boyd, Steven
* Review of tasks and activities identified at last meeting
15 19 Boyd, Steven
** Vacuum / Hydrogen system work : Mark Tucker
16
** Running with Focus Coil : MOM 
17
* Upcoming activities : chaired by DC
18 21 Boyd, Steven
** "This week's MICE Hall work plan":http://micewww.pp.rl.ac.uk/attachments/download/5313  (Paul)
19 18 Boyd, Steven
** Running from Sunday 13th to Thursday 17th 
20 21 Boyd, Steven
*** Is it feasible to run with an absorber? Which absorber? (Chris, All)
21 22 Boyd, Steven
*** If not, what else do we do? FC in solenoid mode is current fallback plan.
22 21 Boyd, Steven
** Personal ODH alarm usage documentation and training (Paul)
23 1 Boyd, Steven
** MOM/DC Handover procedure (Andy)
24 23 Boyd, Steven
** Christmas break walkthrough plans
25 1 Boyd, Steven
26 9 Boyd, Steven
* AOB
27
** [[WorkPlanJanuary2016|To-do list]] for Maintenance shutdown (January 2016)
28
29
h2. Minutes
30 24 Boyd, Steven
31
Present : Paul, Celeste, Sophie, Victoria, Henry, Paolo, Andy, Chris R, Steve W, Steve B, Paul H, Josef, John G, Mark Tucker, Francois,
32
Ken Long, Colin Whyte, Melissa, John C, Yordan, Craig Macwaters, A Gallagher, Ed Overton, P. Hanlet
33
34
Review:
35
36
Vacuum work : most cables are done now. Should be finishing work off this week and then need to get to the stage
37
where controls work. 
38
Hydrogen testing : leak testing done and localised. 
39
Data taking : 4 days of good beam with 1 day of ISIS downtime. Took data for TOF0 alignment. Then data with FC 100A in flip mode
40
and -100 A in flip mode. 
41
42
Plans:
43
44 27 Boyd, Steven
*Tracker He system work : John will work on this. Leak is large and probably can't run with He in the US tracker. Leak is roughly
45 24 Boyd, Steven
4 bottles of He / day and is hard to deal with without taking the patch panel apart. Can we try to gum it up around the seal of
46 25 Boyd, Steven
the butterfly? If we decide to run next week with absorber we need to get He in the tracker volume somehow. 
47 24 Boyd, Steven
48 27 Boyd, Steven
*A Gallagher : Working on vacuum system
49 24 Boyd, Steven
              Taking out the SSU dump trays and taking them back to DL. Americans don't give this a high priority at the moment.
50 1 Boyd, Steven
51 27 Boyd, Steven
*M Courthold : wants to do a 50 A test on the DS whilst it's cold some time this week.*
52
Colin : On that note, DL have contacted FUG with bespoke modifications to the DS PSU and FUG have agreed to the modifications. Hope to have
53 24 Boyd, Steven
a working PS back in February.
54 1 Boyd, Steven
55 27 Boyd, Steven
*Absorber running discussion : this was a baseline plan. We may not be able to run with Xe in the absorber due 
56
to safety and cost concerns. We would have to put 1200 L of Xe in the buffer tank. Andy looked into the 
57
case for Xe - we have to get a permit to do it and doesn't think we can do it in time. Can we work around 
58
the buffer tank? Yes, but it's a significant engineering project. Paul says that the only thing we
59 25 Boyd, Steven
can do in the time available is room temperature nitrogen. John C: Nitrogen at NTP is just as bad as air. A short 
60
material comparison discussion (scat-mat.pdf) showed that only Xe realistically works.
61 24 Boyd, Steven
62 25 Boyd, Steven
A couple of questions arose:
63
Q1. How much is 1200l of Xe and what is the delivery time?
64
Q2. Can we get the safety case signed off and a permit issued before the end of the week. Andy will try.
65
66 1 Boyd, Steven
Current plan : let's try to plan for Xe but expect to do measurements with the FC in solenoid mode.
67 24 Boyd, Steven
68 27 Boyd, Steven
*Christmas walkthrough : Victoria will organise plans for Christmas walkthrough.
69 24 Boyd, Steven
70 27 Boyd, Steven
*AOB: January list: Paul will hunt down system owners and ask them what they want to do in January so that we can develop a plan.
71 24 Boyd, Steven
72
ALH : Spreadsheets need to be filled out and given to Pierrick.
73 25 Boyd, Steven
Henry : need to sort out the temperature monitoring of Q4 and Q5 which have been swapped over after the user period is over. We
74
can do this after the user run finishes on the 18th.
75 24 Boyd, Steven
76 25 Boyd, Steven
John : wants to flush Q5 and Q6 in January. This has to go into the January plan.